# Higher order linear differential equations for unitary matrix integrals: applications and generalisations

### v2

@@ B.4 $D$-finiteness
$y_{l}(z):=\det\left[f_{j-k}(z)\right]_{j,k=0,\ldots,l-1},\quad w_{l}(z):=c_{l}(-1)^{l(l-1)/2}z^{-l^{2}}\det\left[f_{j+k+1}(z)\right]_{j,k=0,\ldots,l-1},$
which are Toeplitz and Hankel determinants of
$f_{n}(z):=I_{n}(2z),$
where $I_{n}$ denotes the modified Bessel function of the first kind of integer order $n$. By the symmetries of the unitary matrix integrals, $y_{l}$ and $w_{l}$ are even entire functions. By choosing the normalization $c_{l}=G(2l+1)/G(l+1)^{2}$ (where $G$ is the Barnes $G$-function), their power series start as [^6] The normalization of $w_{l}$ follows from (2.9) by noting that $_{0}^{\mathstrut}F_{1}^{(1)}(2l;(z^{2})^{l})=1+\frac{z^{2}}{2}+\cdots.$
$y_{l}(z)=1+z^{2}+\cdots,\quad w_{l}(z)=1+\frac{z^{2}}{2}+\cdots.$
Bessel’s differential equation
$z^{2}f_{n}^{\prime\prime}+zf_{n}^{\prime}-(4z^{2}+n^{2})f_{n}=0$
shows that the $f_{n}$ are $D$-finite (holonomic), that is, they satisfy a linear differential equation with polynomial coefficients. Since $D$-finite power series constitute a subalgebra of the formal power series ${\mathbb{C}}\llbracket z\rrbracket$ (see [51, Thm. 2.3]) it is immediately clear that $y_{l},w_{l}$ are also $D$-finite. A symmetry argument shows that the holonomic equations for the even functions $y_{l}$ and $w_{l}$ transform to holonomic equations of the same order for $Y_{l}(x):=y_{l}(\sqrt{x})$ and $W_{l}(x):=w_{l}(\sqrt{x})$ – so these entire functions are also $D$-finite. $D$-finite power series have $P$-recursive (holonomic) coefficients (and vice versa, see [51, Thm. 1.5]), that is, the coefficients satisfy a linear recurrence equation with polynomial coefficients. Besides being of structural interest, such a recurrence allows for the effective calculation of the power series coefficients. Though the general theory of holonomic systems provides simple bounds on the order of the differential equations and recurrences – so that they can be constructed by Groebner bases elimination methods, see [56, Sect. 4] – those bounds are often too large to be of practical use. Since the Bessel differential equation is of order $2$, the degree of the holonomic differential equation for $y_{l}$, or $w_{l}$, is superficially bounded by $2^{l}l!$ because the determinants are sums of $l!$ terms, each of which is the product of $l$ such Bessel functions.
@@
$2zf_{n+1}=zf_{n}^{\prime}-nf_{n},\quad 2z^{2}f_{n+1}^{\prime}=-(n+1)zf_{n}^{\prime}+(4z^{2}+n(n+1))f_{n}.$
@@
$y_{l},w_{l}\in V_{l}:=\text{span}_{{\mathbb{C}}(z)}\left\{(f_{0})^{j}(f_{0}^{\prime})^{l-j}:j=0,\ldots,l\right\};$
@@
the same holds for $w_{l}$. The $a_{j}$ are computed by solving the homogeneous linear system of dimension $(l+1)\times(l+2)$ over ${\mathbb{C}}(z)$ of full rank that is obtained by setting the coefficients of the monomials $f_{0}^{j}(f_{0}^{\prime})^{l-j}$ on the left of (B.1) to zero. Using a computer algebra system, it is thus straightforward to reclaim the differential equations tabulated in Propositions 4.2 and 4.5.
@@
In [10], using the order bound $l+1$, the differential equations, such as (1.6) for $l=4$, were obtained in a far less efficient way by first tabulating, using other means, the first $N$ entries of the power series of $Y_{l}(x)=y_{l}(\sqrt{x})$ for sufficiently large $N$. Using degree bounds for the coefficient polynomials $a_{j}$, the presumed differential equation generates a set of linear equations for the coefficients of $a_{j}$ (normalized to monic $a_{l+1}$) being compatible with the tabulated values. If $N$ is large enough, and the bounds are chosen appropriately, the linear system has a unique solution. This procedure has been implemented in the Maple package gfun [48]; as an example, the differential equation (1.6) is obtained by the command listtodiffeq after adding three more terms to the list (1.2) as follows:
@@
gfun:-Parameters(’maxordereqn’ = l + 1, ’maxdegcoeff’ = l):
@@
B.5 Complexity of tabulating $T_{l}(N)$
As already noted, $T_{l}(N)=N!$ for $N\leq l$, so compiling a table of those numbers of combinatorial interest can be restricted to the triangular array $1\leq l\leq N\leq N_{*}$. Here we compare a tabulation as based on (3.15), to that used in the recent works [29, 12]. [^8] For a history of tabulation by combinatorial methods, see [12, p. 916]. A tabulation for $N_{*}=1000$, computed with the method in [12, §3.2], is available for download at https://box-m3.ma.tum.de/f/7c4f8cb22f5d425f8cff/. There, instead of considering (1.1) directly, one first aims to provide a power series expansion of what is essentially the logarithmic derivative (with $z={1\over 2}\sqrt{s}$),
$u(s):=-s{d\over ds}\log\bigg(e^{-s/4}\Big\langle e^{{1\over 2}\sqrt{s}{\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}\bigg).$ (B.2)
@@ B.5 Complexity of tabulating $T_{l}(N)$
$(su^{\prime\prime})^{2}-(lu^{\prime})^{2}+u^{\prime}(4u^{\prime}-1)(u-su^{\prime})=0$ (B.3)
@@
and substituting in (B.3) provides a recurrence for successive coefficients. However, since the differential equation (B.3) is cubic as a polynomial expression of $u,u^{\prime},u^{\prime\prime}$, the recursion is cubic in the coefficients $c_{n}$, involving double sums for the evaluation of the Cauchy products. It was therefore suggested in [12, §3.2], to use the equivalent third-order first-degree Chazy-I equation, which is obtained from differentiating (B.3) and dividing the result by $2u^{\prime\prime}$:
$s^{2}u^{\prime\prime\prime}+su^{\prime\prime}-6su^{\prime 2}+4uu^{\prime}+(s-l^{2})u^{\prime}-\frac{1}{2}u=0.$ (B.5)
The differential equation is now quadratic as a polynomial expression in $u,u^{\prime},u^{\prime\prime},u^{\prime\prime\prime}$ which yields a much simpler recursive formula for the $c_{n}$, $n=l+1,\ldots$ – quadratic in the $c_{n}$ with just a single sum for the evaluation of the Cauchy products:
$(n+1)(n^{2}-l^{2})c_{n+1}+(n-\tfrac{1}{2})c_{n}-2\sum_{m=l+1}^{n-l}mc_{m}\cdot(3(n-m)+1)c_{n+1-m}=0.$ (B.6)
@@
$O((N_{*}-l)^{2})+O(N_{*}\log N_{*})$
operations. From this, we get the first row in Table B.5.999Note that the total cost when using the $\sigma$-Painlevé III equation as in [29] is $O(N_{*}^{4})$.
@@
  Chazy-I (B.5) $O(N_{*}^{3})$ $O((N_{*}-l)^{2})+O(N_{*}\log N_{*})$ $O(N_{*}^{2})$ $(1-c)^{2}\cdot O(N_{*}^{2})$
recursion (3.15) $O(N_{*}^{3})$ $O(l\cdot N_{*})$ $O(N_{*}^{1+\theta})$ $c\cdot O(N_{*}^{2})$
@@
So, whereas the Chazy-I based computation can take advantage of the shorter rows for $l\approx N_{*}$, a computation based on the recursion (3.15) benefits from the underlying comparatively small dimension $l$ for rows with $l\ll N_{*}$. Overall, when computing the full table, these differences are balanced, and both methods exhibit the same $O(N_{*}^{3})$ complexity.

### v3

@@ B.1 $\boldsymbol{D}$-finiteness
$y_{l}(z):=\det\left[f_{j-k}(z)\right]_{j,k=0,\ldots,l-1},\qquad w_{l}(z):=c_{l}(-1)^{l(l-1)/2}z^{-l^{2}}\det\left[f_{j+k+1}(z)\right]_{j,k=0,\ldots,l-1},$
which are Toeplitz and Hankel determinants of $f_{n}(z):=I_{n}(2z)$, where $I_{n}$ denotes the modified Bessel function of the first kind of integer order $n$. By the symmetries of the unitary matrix integrals, $y_{l}$ and $w_{l}$ are even entire functions. By choosing the normalization $c_{l}=G(2l+1)/G(l+1)^{2}$ (where $G$ is the Barnes $G$-function), their power series start as [^6] The normalization of $w_{l}$ follows from (2.5) by noting that $_{0}^{\mathstrut}F_{1}^{(1)}\bigl(2l;\bigl(z^{2}\bigr)^{l}\bigr)=1+\frac{z^{2}}{2}+\cdots.$
$y_{l}(z)=1+z^{2}+\cdots,\qquad w_{l}(z)=1+\frac{z^{2}}{2}+\cdots.$
Bessel’s differential equation $z^{2}f_{n}^{\prime\prime}+zf_{n}^{\prime}-\bigl(4z^{2}+n^{2}\bigr)f_{n}=0$ shows that the $f_{n}$ are $D$-finite (holonomic), that is, they satisfy a linear differential equation with polynomial coefficients. Since $D$-finite power series constitute a subalgebra of the formal power series ${\mathbb{C}}\llbracket z\rrbracket$ (see [51, Theorem 2.3]), it is immediately clear that $y_{l}$, $w_{l}$ are also $D$-finite. A symmetry argument shows that the holonomic equations for the even functions $y_{l}$ and $w_{l}$ transform to holonomic equations of the same order for $Y_{l}(x):=y_{l}\bigl(\sqrt{x}\bigr)$ and $W_{l}(x):=w_{l}\bigl(\sqrt{x}\bigr)$ – so these entire functions are also $D$-finite. $D$-finite power series have $P$-recursive (holonomic) coefficients (and vice versa, see [51, Theorem 1.5]), that is, the coefficients satisfy a linear recurrence equation with polynomial coefficients. Besides being of structural interest, such a recurrence allows for the effective calculation of the power series coefficients. Though the general theory of holonomic systems provides simple bounds on the order of the differential equations and recurrences – so that they can be constructed by Groebner bases elimination methods, see [56, Section 4] – those bounds are often too large to be of practical use. Since the Bessel differential equation is of order $2$, the degree of the holonomic differential equation for $y_{l}$, or $w_{l}$, is superficially bounded by $2^{l}l!$ because the determinants are sums of $l!$ terms, each of which is the product of $l$ such Bessel functions.
@@
$2zf_{n+1}=zf_{n}^{\prime}-nf_{n},\qquad 2z^{2}f_{n+1}^{\prime}=-(n+1)zf_{n}^{\prime}+\bigl(4z^{2}+n(n+1)\bigr)f_{n}.$
@@
$y_{l},w_{l}\in V_{l}:=\text{span}_{{\mathbb{C}}(z)}\big\{(f_{0})^{j}(f_{0}^{\prime})^{l-j}\colon j=0,\ldots,l\big\};$
@@
the same holds for $w_{l}$. The $a_{j}$ are computed by solving the homogeneous linear system of dimension $(l+1)\times(l+2)$ over ${\mathbb{C}}(z)$ of full rank that is obtained by setting the coefficients of the monomials $f_{0}^{j}(f_{0}^{\prime})^{l-j}$ on the left of (B.1) to zero. Using a computer algebra system, it is thus straightforward to reclaim the differential equations tabulated in Propositions 4.2 and 4.6.
@@
In [11], using the order bound $l+1$, the differential equations, such as (1.6) for l=4{l=4}, were obtained in a far less efficient way by first tabulating, using other means, the first $N$ entries of the power series of $Y_{l}(x)=y_{l}\bigl(\sqrt{x}\bigr)$ for sufficiently large $N$. Using degree bounds for the coefficient polynomials $a_{j}$, the presumed differential equation generates a set of linear equations for the coefficients of $a_{j}$ (normalized to monic $a_{l+1}$) being compatible with the tabulated values. If $N$ is large enough, and the bounds are chosen appropriately, the linear system has a unique solution. This procedure has been implemented in the Maple package gfun [48]; as an example, the differential equation (1.6) is obtained by the command listtodiffeq after adding three more terms to the list (1.2) as follows:
@@
gfun:-Parameters(‘maxordereqn’ = l + 1, ‘maxdegcoeff’ = l):
@@
B.2 Complexity of tabulating $\boldsymbol{T_{l}(N)}$
As already noted, $T_{l}(N)=N!$ for $N\leq l$, so compiling a table of those numbers of combinatorial interest can be restricted to the triangular array $1\leq l\leq N\leq N_{*}$. Here we compare a tabulation as based on (3.12), to that used in the recent works [13, 29]. [^8] For a history of tabulation by combinatorial methods, see [13, p. 916]. A tabulation for $N_{*}=1000$, computed with the method in [13, Section 3.2], is available for download at https://box-m3.ma.tum.de/f/7c4f8cb22f5d425f8cff/. There, instead of considering (1.1) directly, one first aims to provide a power series expansion of what is essentially the logarithmic derivative (with $z={1\over 2}\sqrt{s}$),
$u(s):=-s{{\rm d}\over{\rm d}s}\log\big({\rm e}^{-s/4}\big\langle{\rm e}^{{1\over 2}\sqrt{s}\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)}\big).$ (B.2)
@@ B.2 Complexity of tabulating $\boldsymbol{T_{l}(N)}$
$\bigl(su^{\prime\prime}\bigr)^{2}-\bigl(lu^{\prime}\bigr)^{2}+u^{\prime}\bigl(4u^{\prime}-1\bigr)\bigl(u-su^{\prime}\bigr)=0$ (B.3)
@@
and substituting in (B.3) provides a recurrence for successive coefficients. However, since the differential equation (B.3) is cubic as a polynomial expression of $u$, $^{\prime}, u′′u^{\prime\prime}$, the recursion is cubic in the coefficients $c_{n}$, involving double sums for the evaluation of the Cauchy products. It was therefore suggested in [13, Section 3.2], to use the equivalent third-order first-degree Chazy-I equation, which is obtained from differentiating (B.3) and dividing the result by $2u^{\prime\prime}$
$s^{2}u^{\prime\prime\prime}+su^{\prime\prime}-6su^{\prime 2}+4uu^{\prime}+\bigl(s-l^{2}\bigr)u^{\prime}-\frac{1}{2}u=0.$ (B.5)
The differential equation is now quadratic as a polynomial expression in $u$, $^{\prime}, u′′u^{\prime\prime}$, $u^{\prime\prime\prime}$ which yields a much simpler recursive formula for the $c_{n}$, $n=l+1,\ldots$ – quadratic in the $c_{n}$ with just a single sum for the evaluation of the Cauchy products
$(n+1)\bigl(n^{2}-l^{2}\bigr)c_{n+1}+(n-\tfrac{1}{2})c_{n}-2\sum_{m=l+1}^{n-l}mc_{m}\cdot(3(n-m)+1)c_{n+1-m}=0.$
@@
$O\bigl((N_{*}-l)^{2}\bigr)+O(N_{*}\log N_{*})$
operations. From this, we get the first row in Table B.2.999Note that the total cost when using the $\sigma$-Painlevé III equation as in [29] is $O\bigl(N_{*}^{4}\bigr)$.
@@
  Chazy-I (B.5) $O\bigl(N_{*}^{3}\bigr)$ $O\bigl((N_{*}-l)^{2}\bigr)+O(N_{*}\log N_{*})$ $O\bigl(N_{*}^{2}\bigr)$ $(1-c)^{2}\cdot O\bigl(N_{*}^{2}\bigr)$
recursion (3.12) $O\bigl(N_{*}^{3}\bigr)$ $O(l\cdot N_{*})$ $O\bigl(N_{*}^{1+\theta}\bigr)$ $c\cdot O\bigl(N_{*}^{2}\bigr)$
@@
So, whereas the Chazy-I based computation can take advantage of the shorter rows for $l\approx N_{*}$, a computation based on the recursion (3.12) benefits from the underlying comparatively small dimension $l$ for rows with $l\ll N_{*}$. Overall, when computing the full table, these differences are balanced, and both methods exhibit the same $O\bigl(N_{*}^{3}\bigr)$ complexity.
Acknowledgements
[…]